# Cosmic Recombination in the Presence of Primordial Magnetic Fields

### v2

@@ Abstract
Primordial magnetic fields (PMFs) may explain observations of magnetic fields on extragalactic scales. They are most cleanly constrained by measurements of cosmic microwave background radiation (CMB) anisotropies. Their effects on cosmic recombination may even be at the heart of the resolution of the Hubble tension. We present the most detailed analysis of the effects of PMFs on cosmic recombination to date. To this end we extend the public magneto-hydrodynamic code ENZO with a new cosmic recombination routine, Monte-Carlo simulations of Lyman-$\alpha$ photon transport, and a Compton drag term in the baryon momentum equation. The resulting code allows us, for the first time, to realistically predict the impact of PMFs on the cosmic ionization history and the clumping of baryons during cosmic recombination. Our results identify the importance of mixing of Lyman-$\alpha$ photons between overdense- and underdense- regions for small PMF strength. This mixing speeds up recombination beyond the speed-up due to clumping. For non-helical PMFs with a Batchelor spectrum we find a surprising dependency of results on ultra-violet magnetic modes. The present study shall serve as a theoretical foundation for a future precise comparison of recombination with PMFs to CMB data.
@@ I Introduction
It seems fair to say that the origin of magnetic fields in the present-day Universe is unknown, and further observations and theoretical work is needed. A seemingly clean argument in favor of PMFs would be their detection at very high redshift, before the epoch of structure formation, which is thought to provide the seed fields required by astrophysical generation mechanism [^1] There are also a number of recent studies attempting to constrain approximately scale-invariant PMFs from the enhanced structure formation on small scales Sanati et al. (2020); Adi et al. (2023). An excellent probe could be the accurate observations of the spectrum of, and the anisotropies in the cosmic microwave background radiation (hereafter, CMB), which has already led to precision determination of cosmological parameters in a $\Lambda$CDM model, as well as its approximate confirmation. Indeed many authors have studied the effects of PMFs on the CMB. They considered anisotropic expansion Barrow et al. (1997), spectral $y$ and $\mu$ distortions Jedamzik et al. (2000); Zizzo and Burigana (2005); Kunze and Komatsu (2014), anisotropies due to Alfven and slow magnetic waves Subramanian and Barrow (1998, 2002); Mack et al. (2002); Lewis (2004a); Kahniashvili and Ratra (2005); Chen et al. (2004); Lewis (2004b); Tashiro et al. (2006); Yamazaki et al. (2006); Giovannini (2006); Kahniashvili and Ratra (2007); Giovannini and Kunze (2008); Yamazaki et al. (2010); Paoletti and Finelli (2011); Shaw and Lewis (2012); Kunze (2011); Paoletti and Finelli (2013); Ballardini et al. (2015); Ade et al. (2016); Sutton et al. (2017); Zucca et al. (2017); Minoda et al. (2021), a changing ionization history due to plasma heating Sethi and Subramanian (2005); Kunze and Komatsu (2014, 2015); Ganc and Sloth (2014); Chluba et al. (2015); Ade et al. (2016), extra polarization due to PMFs Durrer et al. (2000); Seshadri and Subramanian (2001); Mack et al. (2002); Subramanian et al. (2003); Mollerach et al. (2004); Lewis (2004a); Scoccola et al. (2004); Kosowsky et al. (2005); Kahniashvili and Ratra (2005); Pogosian et al. (2013); Kahniashvili et al. (2014); Zucca et al. (2017); Pogosian and Zucca (2018), as well as non-Gaussianity in bi- and tri-spectrum  Brown and Crittenden (2005); Seshadri and Subramanian (2009); Caprini et al. (2009); Cai et al. (2010); Trivedi et al. (2010); Brown (2011); Shiraishi et al. (2010, 2011); Trivedi et al. (2012); Shiraishi and Sekiguchi (2014); Trivedi et al. (2014); Ade et al. (2016). Most works have not seen evidence for PMFs and have imposed upper limits in the comoving $\sim$ nG range [^2] The exception here is the limit on PMFs of $B_{0}<0.05nG$ Trivedi et al. (2014) on inflationary fields from the non-Gaussianity in the tri-spectrum when the inflationary curvature mode is taken into account.. It is noted that these upper limits are much larger than the field strength $B_{0}\sim 0.005\,$nG Banerjee and Jedamzik (2003). required for clusters of galaxies to have an exclusively primordial origin.
@@
The outline of the paper is as follows. In Section II we discuss details of the numerical MHD simulations, the employed MHD solver, the physical effects included, and the newly written recombination routine. We then present results of a realistic baryon pdf. Section III reviews standard recombination theory and presents results of a Monte-Carlo analysis of the propagation of Lyman-alpha photons during recombination. In Section IV we analyze the effects of Lyman-$\alpha$ photon mixing on the ionization history, presenting a new explicit expression for the recombination rate in the full-mixing limit. Section V presents results of MHD simulations of PMFs before, during, and after recombination. Particular emphasis is laid on non-helical PMFs with a Batchelor spectrum and the change of the global free electron fraction $X_{e}\equiv\langle{n_{e}}\rangle/\langle{n_{\rm H}}\rangle$ (i.e. the ratio of the average free-electron abdundance to the average abundance of hydrogen nuclei) compared to a non-magnetized Universe. Conclusions are drawn in Section VI. In Appendix A we present the details of our Lyman-$\alpha$ Monte-Carlo simulation. Appendix B briefly describes helium recombination. Appendix C presents a numerical convergence study.
@@ III.2 Lyman-$\alpha$ mixing length from Monte Carlo simulations
Fig. 5 investigates the fraction of Lyman-$\alpha$ photons which is lost due to redshifting, as a function of bulk flows. Without bulk flows, in a Universe with cosmological parameters as inferred from the best-fit to the Planck data, the fraction of photons lost due to redshifting and lost due to two photon decay are $f_{z}({\rm v}=0)\approx 5\times 10^{-3}$ and $f_{2\gamma}\approx 1.5\times 10^{-2}$ at redshift $z\approx 1100$, respectively. In the presence of peculiar flows results are very dependent of the coherence scale $\lambda$ of the flow. The yellow line shows result for $\lambda\gg D_{\alpha}$ and is in full agreement with the analytical result. However, for smaller $\lambda$ the effect is much diminished as the Lyman-alpha photons alternatively travel through diverging and converging flows. For $\lambda=0.1\,$kpc the result of $f_{z}({\rm v}=0)$ is essentially recovered, whereas for $\lambda=1\,$kpc there is a small deviation from $f_{z}({\rm v}=0)$, with however a reduction for ${\bf\nabla\cdot v}<0$ and an enhancement for ${\bf\nabla\cdot v}>0$. In Section V we perform MHD simulations for a variety of magnetic field strength. From these simulations we can infer at typical $\lambda\approx 0.1\,$kpc and 111\,1kpc for PMFs with a Batchelor spectrum leading to a ”final” field strength of 202020\,20pG and 136136136\,136pG, respectively. We conclude therefore that for low final field strengths the effects of peculiar flows on the Lyman-$\alpha$ escape rate should be negligible, whereas for much stronger PMFs, we expect an additional reduction of $X_{e}$ due to peculiar flows. However, these latter field strength seem already ruled out if only the effect of clumping is taken into account Jedamzik and Saveliev (2019).
@@ V.2 Optimizing calculations for non-helical magnetic fields with a Batchelor spectrum
VI Conclusions

### v3

@@ Abstract
Primordial magnetic fields (PMFs) may explain observations of magnetic fields on extragalactic scales. They are most cleanly constrained by measurements of cosmic microwave background radiation (CMB) anisotropies. Their effects on cosmic recombination may even be at the heart of the resolution of the Hubble tension. We present the most detailed analysis of the effects of PMFs on cosmic recombination to date. To this end we extend the public magneto-hydrodynamic code ENZO with a new cosmic recombination routine, Monte-Carlo simulations of Lyman-$\alpha$ photon transport, and a Compton drag term in the baryon momentum equation. The resulting code allows us, for the first time, to realistically predict the impact of PMFs on the cosmic ionization history and the clumping of baryons during cosmic recombination. Our results identify the importance of mixing of Lyman-$\alpha$ photons between overdense- and underdense- regions for small PMF strength. This mixing speeds up recombination beyond the speed-up due to clumping. We also investigate the effects of pecuilar flows on the recombination rate and find it to be small for small PMF strengths. For non-helical PMFs with a Batchelor spectrum we find a surprising dependency of results on ultra-violet magnetic modes. We further show that the increase in the ionization fraction at low redshift by hydrodynamic baryon heating due to PMF dissipation is completely compensated by the faster recombination from baryon clumping. The present study shall serve as a theoretical foundation for a future precise comparison of recombination with PMFs to CMB data.
@@ I Introduction
It seems fair to say that the origin of magnetic fields in the present-day Universe is unknown, and further observations and theoretical work is needed. A seemingly clean argument in favor of PMFs would be their detection at very high redshift, before the epoch of structure formation, which is thought to provide the seed fields required by astrophysical generation mechanism [^1] There are also a number of recent studies attempting to constrain approximately scale-invariant PMFs from the enhanced structure formation on small scales Sanati et al. (2020); Adi et al. (2023). An excellent probe could be the accurate observations of the spectrum of, and the anisotropies in the cosmic microwave background radiation (hereafter, CMB), which has already led to precision determination of cosmological parameters in a $\Lambda$CDM model, as well as its approximate confirmation. Indeed many authors have studied the effects of PMFs on the CMB. They considered anisotropic expansion Barrow et al. (1997), spectral $y$ and $\mu$ distortions Jedamzik et al. (2000); Zizzo and Burigana (2005); Kunze and Komatsu (2014), anisotropies due to Alfven and slow magnetic waves Subramanian and Barrow (1998, 2002); Mack et al. (2002); Lewis (2004a); Kahniashvili and Ratra (2005); Chen et al. (2004); Lewis (2004b); Tashiro et al. (2006); Yamazaki et al. (2006); Giovannini (2006); Kahniashvili and Ratra (2007); Giovannini and Kunze (2008); Yamazaki et al. (2010); Paoletti and Finelli (2011); Shaw and Lewis (2012); Kunze (2011); Paoletti and Finelli (2013); Ballardini et al. (2015); Ade et al. (2016a); Sutton et al. (2017); Zucca et al. (2017); Minoda et al. (2021), a changing ionization history due to plasma heating Sethi and Subramanian (2005); Kunze and Komatsu (2014, 2015); Ganc and Sloth (2014); Chluba et al. (2015); Ade et al. (2016a), extra polarization due to PMFs Durrer et al. (2000); Seshadri and Subramanian (2001); Mack et al. (2002); Subramanian et al. (2003); Mollerach et al. (2004); Lewis (2004a); Scoccola et al. (2004); Kosowsky et al. (2005); Kahniashvili and Ratra (2005); Pogosian et al. (2013); Kahniashvili et al. (2014); Zucca et al. (2017); Pogosian and Zucca (2018), as well as non-Gaussianity in bi- and tri-spectrum  Brown and Crittenden (2005); Seshadri and Subramanian (2009); Caprini et al. (2009); Cai et al. (2010); Trivedi et al. (2010); Brown (2011); Shiraishi et al. (2010, 2011); Trivedi et al. (2012); Shiraishi and Sekiguchi (2014); Trivedi et al. (2014); Ade et al. (2016a). Most works have not seen evidence for PMFs and have imposed upper limits in the comoving $\sim$ nG range [^2] The exception here is the limit on PMFs of $B_{0}<0.05nG$ Trivedi et al. (2014) on inflationary fields from the non-Gaussianity in the tri-spectrum when the inflationary curvature mode is taken into account.. It is noted that these upper limits are much larger than the field strength $B_{0}\sim 0.005\,$nG Banerjee and Jedamzik (2003). required for clusters of galaxies to have an exclusively primordial origin.
@@
The outline of the paper is as follows. In Section II we discuss details of the numerical MHD simulations, the employed MHD solver, the physical effects included, and the newly written recombination routine. We then present results of a realistic baryon pdf. Section III reviews standard recombination theory and presents results of a Monte-Carlo analysis of the propagation of Lyman-alpha photons during recombination. In Section IV we analyze the effects of Lyman-$\alpha$ photon mixing on the ionization history, presenting a new explicit expression for the recombination rate in the full-mixing limit. Section V presents results of MHD simulations of PMFs before, during, and after recombination. Particular emphasis is laid on non-helical PMFs with a Batchelor spectrum and the change of the global free electron fraction $X_{e}\equiv\langle{n_{e}}\rangle/\langle{n_{\rm H}}\rangle$ (i.e. the ratio of the average free-electron abdundance to the average abundance of hydrogen nuclei) compared to a non-magnetized Universe. In Section VI we discuss the combined effect of baryon clumping and hydrodynamic heating by PMF dissipation on $X_{e}$. Conclusions are drawn in Section VII. In Appendix A we present the details of our Lyman-$\alpha$ Monte-Carlo simulation. Appendix B briefly describes helium recombination. Appendix C presents a numerical convergence study.
@@ II.1 Physical problem to be solved
Though the baryon temperature $T_{b}$ is to excellent approximation given by the CMB temperature $T_{\rm CMB}$ due to Thomson scattering for redshifts $z\gtrsim 1000$, it may deviate at lower redshifts. In particular, $T_{b}$ may exceed $T_{\rm CMB}$ due to the energy released into the plasma by magnetic field dissipation Sethi and Subramanian (2005). The evolution of the local baryon temperature is governed by the following equation
$\frac{dT_{b}}{dt}=\frac{2}{3}\frac{d\rho_{b}/dt}{\rho_{b}}+\frac{8\sigma_{Th}x_{e}n_{H}\rho_{\gamma}}{3m_{e}cn_{\rm tot}}\bigl{(}T_{\rm CMB}-T_{b})+\frac{2\Gamma}{3n_{\rm tot}}\,,$ (8)
where the the terms on the RHS are due to adiabatic cooling/heating, CMB cooling/heating, and heating due to magnetic field dissipation, with $\Gamma$ the volumetric heating rate. We will discuss $\Gamma$ and the effects of this dissipation in Section VI.
@@ III.2 Lyman-$\alpha$ mixing length from Monte Carlo simulations
Fig. 5 investigates the fraction of Lyman-$\alpha$ photons which is lost due to redshifting, as a function of bulk flows. Without bulk flows, in a Universe with cosmological parameters as inferred from the best-fit to the Planck data, the fraction of photons lost due to redshifting and lost due to two photon decay are $f_{z}({\rm v}=0)\approx 5\times 10^{-3}$ and $f_{2\gamma}\approx 1.5\times 10^{-2}$ at redshift $z\approx 1100$, respectively. In the presence of peculiar flows results are very dependent of the coherence scale $\lambda$ of the flow. The yellow line shows result for $\lambda\gg D_{\alpha}$ and is in full agreement with the analytical result. However, for smaller $\lambda$ the effect is much diminished as the Lyman-alpha photons alternatively travel through diverging and converging flows. For $\lambda=0.1\,$kpc the result of $f_{z}({\rm v}=0)$ is essentially recovered, whereas for $\lambda=1\,$kpc there is a small deviation from $f_{z}({\rm v}=0)$, with however a reduction for ${\bf\nabla\cdot v}<0$ and an enhancement for ${\bf\nabla\cdot v}>0$. In Section V we perform MHD simulations for a variety of magnetic field strength. From these simulations we can infer at typical $\lambda\approx 0.1\,$kpc and 111\,1kpc for PMFs with a Batchelor spectrum leading to a ”final” field strength of 202020\,20pG and 136136136\,136pG, respectively. We suspect therefore that for low final field strengths the effects of peculiar flows on the Lyman-$\alpha$ escape rate should be negligible, whereas for much stronger PMFs, we expect an additional reduction of $X_{e}$ due to peculiar flows. However, this latter field strength seems already ruled out if only the effect of clumping is taken into account Jedamzik and Saveliev (2019).
It is our goal to obtain a very precise prediction of the average ionzation fraction $X_{e}$. To verify our suspicion that additional losses of Lyman-$\alpha$ photons due to bulk flows are negligible, we have indulged in a more elaborate Monte Carlo simulation of Lyman-$\alpha$ photon loss. In particular, we have used snapshots of the numerical simulations in order to have a realistic and detailed knowledge of the density- and velocity- distribution at a given redshift. We have then ”emitted” Lyman-$\alpha$ photons from a great number of randomly chosen locations and followed their propagation in physical space and evolution of energy space. This allowed us to compute the global Lyman-$\alpha$ photon loss rate numerically   [^8] In particular we evaluated the integral $(1/V)\int C\alpha_{e}n_{e}^{2}$ by Monte Carlo integration and compared it to the fully Lyman-$\alpha$ mixed (cf. Section IV) result $(1/V)\,C^{\rm mix}\int\alpha_{e}n_{e}^{2}$ which was applied in the simulations. and compare it to the one in a Universe without pecuilar bulk flows and fully mixed Lyman-$\alpha$ photons. Differences between these two global loss rates are due to two effects: bulk flows and the effect of partial Lyman-$\alpha$ mixing important at higher redshifts where $D_{\alpha}$ becomes relatively small. To isolate the two effects we ran our Monte-Carlo simulations with and without pecuilar flows. For a simulation with Batchelor spectrum leading to a final field of 20 pG we did find an enhancement of the loss reate by only a few percent, with the loss rate at the most important redshifts around recombination $z\sim 1100$ virtually identical. The enhancement of a few per cent was in almost equal parts due to bulk flows and partial Lyman-$\alpha$ photon mixing. By computing the changes in ionization history due to the small enhancement of Lyman-$\alpha$ photon loss in a homogeneous Universe (for simplicity) we could conclude that the relative change should be below the one per cent level for those field strengths not already ruled out. We conclude therefore that negelcting bulk flows and assuming full Lyman-$\alpha$ mixing introduces errors in $X_{e}$ not larger than one per cent.
@@ V.2 Optimizing calculations for non-helical magnetic fields with a Batchelor spectrum
VI The combined effects of plasma heating and baryon clumping
It is well known Sethi and Subramanian (2005) that the decay of magnetic turbulence induces a heat source for the baryons which may influence the ionization fraction $X_{e}$ substantially at lower redhifts. Similar holds for ambipolar diffusion at even lower redshifts. An increase of the baryon temperature $T_{b}$ over that of the CMB $T_{\rm CMB}$ lowers the temperature-dependant recombination rate which leads to higher residual ionization fractions. The evolution of the baryon temperature with time $t$ is given by (8). At redshifts $z\approx 1000$ the ionization fraction is still fairly high $X_{e}\approx 0.1$ such that baryon cooling by electron Thomson scattering on CMB photons keeps $T_{b}$ very close to $T_{\rm CMB}$. However, at slightly lower redshifts $z\approx 900$ when $X_{e}\approx 10^{-2}$ the heating of baryons due to the dissipation of magnetic fields can significantly increase the baryon temperature. This in turn reduced the temperature dependent recombination rate which leads to higher residual $X_{e}$.
Figure 12: The ionization fraction $X_{e}$ as a function of redshift for four different cases: the homogeneous case without clumping and dissipation (orange), the inhomogeneous case due to baryon clumping, but without dissipation (blue), the homogeneous case without clumping but with dissipation (purple), and the realistic inhomogeneous case with clumping and dissipation (green).
A higher residual $X_{e}$ in the redshift range $\sim 400-900$ induces a suppression of the CMB anisotropies at higher multipoles due to residual scatterings of photons out of the beam. It also introduces a feature at small multipoles $l\sim 100-200$. These effects may be constrained by high-precision CMB data, as has been done in Kunze and Komatsu (2015); Chluba et al. (2015); Ade et al. (2016b); Paoletti et al. (2019, 2022). These authors used an analytic expression from Sethi and Subramanian (2005) to approximate the heating rate due to magnetic field dissipation. The only study which attemps to compute the heating rate numerically Trivedi et al. (2018) could not simulate at higher magnetic field strength, as they were facing the complication of baryon clumping, which many MHD solvers have problems with. We introduced heating of baryons due to magnetic field dissipation in the modified ENZO code. To this end we use a uniform heating rate   [^11] Magnetic- and kinetic- energy dissipation here occurs over numerical viscosity, which is a common practice in MHD simulations in astrophysics Kritsuk et al. (2011). In principle we could have computed the local dissipation rate when using the standard energy evolution equation in ENZO amended by CMB cooling/heating, as ENZO conserves energy. However, we do not think that the approximation of applying a global average homogeneous heating rate could change our conclusions below drastically.
[…]
which allows us to evolve the local baryon temperature. We did not consider the less important heating due to ambipolar diffusion. We are thus able, for the first time, to treat both effects, clumping and heating at the same time.
Results are shown in Fig. 12. Here a particularly strong magnetic field with Batchelor spectrum was chosen, with remaining magnetic field strength of 1.21.21.2\,1.2nG at $z=1088$ leading to a final field of 0.130.130.13\,0.13nG at redshift $z=10$. According to Paoletti et al. (2022), which give the limit $B<0.18\,$nG at $z=1088$ such a field should be clearly ruled out due to heating. It should also be ruled out due to clumping, if the three-zone model M1 with older data is used Jedamzik and Saveliev (2019), due to modifications of $X_{e}$ at $z\approx 1100$ rather than at $z\approx 400-900$. In Fig. 12 one sees the ionization fraction for four different simulations. The orange line shows $X_{e}$ for a homogeneous Universe with no magnetic fields. The purple line shows $X_{e}$ when the heating rate is extracted from the MHD simulation and used in a homogeneous Universe. This corresponds to the evolution of $X_{e}$ computed in Paoletti et al. (2022) with the difference that the heating rate is not taken from an analytic approximation but computed directly in the code. This case seems clearly ruled out by current CMB data. The blue line shows $X_{e}$ when heating is not taken into account, but the baryon inhomogeneities, i.e. clumping, are. In this case the average $X_{e}$ is lower than in the no magnetic field case, and taking only the low $z$ evolution into account the model should not be ruled out. The same holds for the realistic case, when clumping and heating is taken into account, which is shown by the green line. The realistic case in fact is very close to the no magnetic field case at low redshifts. Baryon heating due to magnetic field dissipation increases $X_{e}$ at lower redshift, whereas the clumping of baryons decreases $X_{e}$ with the net effect being only small change from the no magnetic field case. In fact we have performed many more simulations with varying magnetic field strengths including the effects of clumping and heating and in none of these simulations have we ever found an increased $X_{e}$. We conclude therefore, that current limits on PMFs from baryon heating do not apply, as they have not taken into account clumping. We stress, that constraints on this model come from clumping, and the associated changes of $X_{e}$ around the peak of the visibility function at $z\approx 1090$, inducing a change in the sound horizon, modified Silk damping, and a change in the width of the visibility function, but not from hydrodynamic heating.
VII Conclusions